Given the following function:

F (X) = -93 - 31X2 - 19X

1. What is the value of the critical point for this function?

2. Is this critical point a maximum, minimum or is it indeterminate?
1
Expert's answer
2019-11-18T14:01:51-0500

(1) To obtain critical value(s), the equation is differentiated and equated to zero. Given "F (X) = -93 - 31X^2 - 19X". The first derivative is given as

"\frac{dy}{dx}(−31x^2−19x−93)"

"=−31\frac{d}{dx}(x^2)−19\frac{d}{dx}(x)+\frac{d}{dx}(−93)"

"=\u221262x\u221219+0=\u221262x\u221219"

Equating to zero, we have

"0=\u221262x\u221219"

"19=-62x"

"x=\frac{-19}{62}"

Then "x" is substituted into "F(X)" to get "y" coordinate. Thus,

"F (\frac{-19}{62}) = -93 - 31(\frac{-19}{62})^2 - 19\frac{-19}{62}"

"=-93-31(\frac{361}{3844})+\frac{361}{62}"

"y=-93-\frac{11191}{3844}+\frac{361}{62}"

"y=-93+\frac{11191}{3844}=-90.0089".

Therefore, the critical point of the equation is at "x=\frac{-19}{62}" and the value "y=-90.0089"

(2). To determine whether the critical point is maximum, minimum or indeterminate, the second derivative is required. The second derivative is given by

"\frac{d}{dx}(−62x−19)"

"=−62\frac{d}{dx}(x)+\frac{d}{dx}(−19)"

"=0\u221262\n=\u221262"

Since the value of the second derivative is negative, the critical point is a maximum.